OpenStudy (anonymous):

Please help me solve. 1) 14 = d/6 -7 2) 9 = x-5/4

pooja195 (pooja195):

1) add 7 to both sides 14+7 = d/6 -7+7 21=d/6 then Multiply by the reciprocal so flip d/6 to 6/d 21(6)=d <---the 6 cancels out so d=21(6) <---figure out what that equals.

pooja195 (pooja195):

2) Simply the equation so that it looks like this 9=x+-5/4 then add -5/4 to both sides 9+-5/4=x+-5/4 +-5/4 x=-5/4+9 <=figure out what it equals

OpenStudy (anonymous):

the second one is right

OpenStudy (anonymous):

thanks

pooja195 (pooja195):

np

pooja195 (pooja195):

\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)